# WAM-Flow: Parallel Coarse-to-Fine Motion Planning via Discrete Flow Matching for Autonomous Driving

### v1

@@ 1 Introduction
Vision-language–action models for end-to-end autonomous driving [62, 58, 32] aim to map egocentric driving-view video inputs and natural-language instructions into both causal reasoning and precise ego-vehicle motion planning, while satisfying stringent efficiency and safety requirements. A fundamental challenge in this domain is the design of a policy representation that effectively balances three critical aspects: expressive reasoning capabilities, high-fidelity continuous control, and robust closed-loop performance. Existing approaches can be broadly categorized into dual-system and single-system paradigms. Dual-system methods [32, 60, 61, 50, 20] typically employ autoregressive vision-language models (VLMs) [46, 36, 48, 3, 63] as auxiliary reasoning modules to provide high-level driving intent, scene summaries, or linguistic guidance for downstream motion planning networks, which often utilize diffusion-based iterative optimization [19, 34, 24, 13] to generate smooth, complex action distributions. In contrast, single-system approaches [23, 52, 7, 58, 62] such as EMMA [23] and DrivingGPT [7] reformulate trajectory or action prediction as a text generation problem within the VLM, enabling reasoning and planning directly in the linguistic space. This work investigates a novel alternative based on discrete flow matching (DFM), which offers distinct advantages for autonomous driving applications.
@@
Discrete flow matching [14, 42, 47, 12, 44, 27, 57, 8] models probability transport over discrete token spaces via a continuous-time Markov chain (CTMC) that carries a simple base distribution to the data distribution. Unlike autoregressive decoders that commit to tokens sequentially and accumulate exposure-bias errors, Discrete flow matching supports fully parallel denoising and bidirectional refinement during generation. These properties enable coarse-to-fine planning: beginning with a coarse motion hypothesis, the model increases trajectory fidelity through additional denoising steps, yielding a tunable compute–accuracy trade-off. This flexibility aligns well with autonomous driving, where simple scenes admit rapid approximate plans while complex interactions require higher-precision refinement. Despite these advantages, discrete flow matching remains largely unexplored for VLA policies in end-to-end autonomous driving.
@@
However, a straightforward application of discrete flow matching to VLA model for end-to-end autonomous driving is nontrivial for three reasons. First, training discrete flow matching from scratch is prohibitively data- and compute-intensive, so they are typically initialized from general-purpose autoregressive multimodal VLMs that lack sufficient road-scene competence–from low-level perception and motion forecasting to high-level planning and decision making. We therefore adopt a multi-stage adaptation strategy: starting from a generic VLM backbone (Janus-1.5B [51]), we continued conduct pretraining on large-scale road-scene visual question answering (VQA) to strengthen the ability to understand various complex road scenes and vehicle driving patterns, establishing a strong domain prior comparable to autoregressive VLA baselines. Second, standard text token embeddings are ill-suited to high-precision numerical regression because they weakly encode metric relationships. We introduce a metric-aligned numerical tokenizer that discretizes continuous scalars into a shared codebook and learns embeddings with a triplet-margin ranking objective so that latent distances reflect underlying scalar differences. This structured token space enables stable coarse-to-fine and slow–fast trajectory refinement within discrete flow matching, providing a controllable compute–accuracy trade-off. Finally, supervised likelihood-based flow training aligns the model with expert trajectories but does not explicitly enforce safety, ego-progress, and comfort in closed-loop control. We incorporate a Group Relative Policy Optimization (GRPO) based alignment objective with a composite reward that integrates safety penalties and performance goals, improving the safety–progress–comfort profile while preserving the model’s parallel generation capabilities.
@@ 2 Related Work
VLMs in Autonomous Driving. Autoregressive VLMs  [23, 52, 58, 62, 45, 25] formulate driving as a sequential language modeling problem, where each token corresponds to a trajectory point, control command, or reasoning step. Representative works such as EMMA [23], OpenEMMA [52], FutureSightDrive [58] and AutoVLA [62] leverage chain-of-thought reasoning and external memory modules to enhance interpretability and decision transparency. Despite their strong causal modeling capability, autoregressive architectures suffer from slow autoregressive decoding and limited parallelism, as future actions must be generated step-by-step. Diffusion-based methods, including DiffusionDrive [34], ViLaD [10] and DiffVLA [24] treat planning as a denoising process that gradually refines latent trajectory representations. These models enable parallel sampling but often lack explicit reasoning interpretability. In this paper, we explore a new promising paradigm for end-to-end autonomous driving, namely discrete flow matching.
@@
Discrete Diffusion in LLMs and VLMs. Recent progress in discrete generative modeling has led to the emergence of discrete diffusion LLMs [39, 54] and discrete diffusion VLMs [38, 56, 55, 30, 53], which extend diffusion processes to tokenized sequences. This direction originates from D3PM [2], which formulated diffusion as a discrete Markov process over categorical variables. Recently, LLaDA [39] trains an 8B-parameter model from scratch, reaching LLaMA-3 [16] performance with bidirectional reasoning and robustness. DREAM-7B [54] further enhances diffusion-based reasoning via iterative refinement and arbitrary-order generation. Meanwhile, discrete flow matching [15, 35, 41] generalizes discrete diffusion via continuous-time probability paths and learnable velocity fields. By unifying diffusion and flow-based generation under a single probabilistic framework, it enables parallel, bidirectional, and efficient sampling. FUDOKI [47] extends this framework to multimodal reasoning and generation, demonstrating unified, non-autoregressive modeling across modalities. In this paper, we apply discrete flow matching to VLA for autonomous driving and explore its inherent nature of parallel generation and coarse-to-fine controllability.
@@
Reinforcement Learning in VLA. Building upon the success of DeepSeek-R1 [17], GRPO has been further extended to autonomous driving domains. In particular, AlphaDrive [26] pioneers the integration of GRPO-based reinforcement learning with planning-centric reasoning in autonomous driving, achieving notable improvements in both decision-making performance and training efficiency. TrajHF [28] further combines diffusion-based multimodal planners with reinforcement learning from human feedback, enabling safe and personalized trajectory generation aligned with diverse human driving styles. More recently, AutoVLA [62] incorporates GRPO into vision-language-action models, extending reinforcement learning to end-to-end multimodal reasoning and low-level planning. To the best of our knowledge, this work presents the first exploration of GRPO within discrete flow matching for autonomous driving VLA. Furthermore, we explicitly incorporate safety alignment objectives, extending beyond conventional likelihood-based training to enhance reliability in autonomous driving contexts.
@@ 3.2 WAM-Flow Architecture
Model Architecture. We adapt a Janus-1.5B multimodal backbone to the discrete flow matching generation paradigm for vision–language–action planning. Images are resized with preserved aspect ratio, zero-padded to 384×384, and encoded by SigLIP [59] into 576 visual tokens; a lightweight MLP aligns these features to the 2048-dimensional Janus text-token space. On the language side, we extend the Janus tokenizer by 20,001 numerically grounded tokens to represent input ego-state numbers and output waypoint coordinates, yielding a 122,401-word vocabulary. Training data are formatted with a fixed QA-style prompt that integrates navigation commands, ego-state (position, heading, velocity, acceleration), and the target waypoint sequence for the next 4 seconds (8 waypoints). For the decoder, the original Janus text head is expanded to the enlarged vocabulary and used to predict action tokens under the discrete flow matching objective.
@@ 3.4 Training and Inference
Autoregressive-to-Flow Training. Figure 2 outlines a four-stage curriculum. First, we randomly initialize the numerical embeddings and freeze the VLA backbone. We train the numerical embeddings together with the language-model head on the 668K nuPlan dataset for 4 epochs, using flow-matching loss $\mathcal{L}_{\mathrm{CE}}$ (Equation 8) and triplet-margin ranking loss $\mathcal{L}_{\mathrm{num}}$ (Equation 5). Second, we enhance the perception of driving scenes by pretraining VLA using $\mathcal{L}_{\mathrm{CE}}$. This stage trains 3 epochs on 6.5M VQA, including general multimodal VQA (3.4M) from LLaVA-v1.5 [36] and large-scale driving-specific VQA (3.1M) from RecogDrive [32], which enhances perceptual grounding and driving-specific causal reasoning. Third, we supervised fine-tuning of the VLA backbone for only 2 epochs on the nuPlan dataset with $\mathcal{L}_{\mathrm{CE}}$. After supervised flow training, we perform reinforcement learning with simulator feedback by maximizing the GRPO objective (Equation 10) with KL regularization toward the supervised reference to optimize our VLA model for 0.5 epoch on 103k NAVSIM dataset. We set the weight for $\mathrm{EP},\mathrm{TTC}$ and $\mathrm{C}$ in our reward to 5:5:2.
@@ 4.1 Experimental Setup
Metrics. We evaluate our method on the closed-loop NAVSIM-v1 [11] and v2 [5] benchmarks. The primary metric for NAVSIM-v1 is the Predictive Driver Model Score (PDMS), a composite measure integrating five key components: No-Collision rate (NC), Drivable Area Compliance (DAC), Time-to-Collision within bound (TTC), Comfort, and Ego Progress (EP). The more comprehensive NAVSIM-v2 benchmark employs the Extended Predictive Driver Model Score (EPDMS), which incorporates nine sub-metrics—NC, DAC, Driving Direction Compliance (DDC), Traffic Light Compliance (TLC), EP, TTC, Lane Keeping (LK), History Comfort (HC), and Extended Comfort (EC)– to provide a holistic assessment of driving performance, safety, and rule adherence. All results are obtained from closed-loop simulations on the official public test splits.
@@ 5 Conclusion
This appendix provides additional experimental results and implementation details to complement the main paper. Specifically, Section A presents extended evaluations on the nuScenes [4] datasets, along with additional qualitative experiments on NAVSIM. Section B provides pseudocode for the training and inference stages, respectively. Section C elaborates on the evaluation metrics, and Section D discusses implementation specifics. Finally, Section E discuss the limitation and future work.
@@ Appendix A Additional Experiments
We evaluate our method on the nuScenes dataset [4] following the NAVSIM benchmark perspective [11, 5], which focuses on collision rate as the primary metric. This emphasis stems from the established finding in NAVSIM that open-loop L2 distance exhibits negligible correlation with closed-loop performance. As shown in Table 7, our method achieves an average collision rate of 0.12% under ST-P3 metrics, matching the performance of the best non-VLA model (UniAD). More notably, under the more comprehensive UniAD metrics, WAM-Flow sets a new state-of-the-art with the lowest average collision rate (0.23%) among all evaluated VLA methods. The model also demonstrates superior short-term safety, achieving a perfect 0.00% collision rate at the 1-second horizon.
@@
PreWorld [31] - - - - - - 0.19 0.57 2.65 1.14
ST-P3 [21] - - 0.23 0.62 1.27 0.71 - - - -
Ego-MLP [33] - - 0.21 0.35 0.58 0.38 - - - -
InsightDrive [43] - - 0.09 0.10 0.27 0.15 0.08 0.15 0.84 0.36
VAD-v2 [6] - - 0.07 0.10 0.24 0.14 - - - -
UniAD [22] - - 0.04 0.08 0.23 0.12 0.05 0.17 0.71 0.31
@@
Epona [60] AR + Diff. DiT-2.5B [40] 0.05 0.22 0.85 0.96 - - - -
OmniDrive [49] AR LLaVA-7B [36] 0.04 0.46 2.32 0.94 - - - -
DriveVLM [45] AR Qwen2-VL-7B [48] 0.10 0.22 0.45 0.27 - - - -
GPT-Driver [37] AR GPT-4 [1] 0.04 0.12 0.36 0.17 0.07 0.15 1.10 0.44
AutoVLA [62] AR Qwen2.5-3B [3] 0.13 0.18 0.28 0.20 0.14 0.25 0.53 0.31
DME-Driver [18] AR LLaVA-7B [36] - - - - 0.05 0.28 0.55 0.29
Ours DFM Janus-1.5B [51] 0.04 0.10 0.23 0.12 0.00 0.10 0.60 0.23
Table 7: End-to-end motion planning performance on the nuScenes [4] dataset. We sort previous methods according to the average collision rate. Abbreviation: Diff.(Diffusion), AR (autoregressive), DFM (discrete flow matching).

### v2

@@ 1 Introduction
Vision-language–action models for end-to-end autonomous driving [zhou2025autovla, FutureSightDrive-2025, RecogDrive-2025] aim to map egocentric driving-view video inputs and natural-language instructions into both causal reasoning and precise ego-vehicle motion planning, while satisfying stringent efficiency and safety requirements. A fundamental challenge in this domain is the design of a policy representation that effectively balances three critical aspects: expressive reasoning capabilities, high-fidelity continuous control, and robust closed-loop performance. Existing approaches can be broadly categorized into dual-system and single-system paradigms. Dual-system methods [RecogDrive-2025, Epona-2025, AdaDrive-2025, Alpamayo-R1-2025, DriveAgent-2025] typically employ autoregressive vision-language models (VLMs) [LLaMA-2, LLaVA, wang2024qwen2, Qwen-25, InternVL3-2025] as auxiliary reasoning modules to provide high-level driving intent, scene summaries, or linguistic guidance for downstream motion planning networks, which often utilize diffusion-based iterative optimization [DDPM-2020, DiffusionDrive-2025, jiang2025diffvla, Artemis-2025] to generate smooth, complex action distributions. In contrast, single-system approaches [EMMA-2024, OpenEMMA-2025, DrivingGPT-2025, FutureSightDrive-2025, zhou2025autovla] such as EMMA [EMMA-2024] and DrivingGPT [DrivingGPT-2025] reformulate trajectory or action prediction as a text generation problem within the VLM, enabling reasoning and planning directly in the linguistic space. This work investigates a novel alternative based on discrete flow matching (DFM), which offers distinct advantages for autonomous driving applications.
@@
Discrete flow matching [DFM-2024, DFM-2025, wang2025fudoki, deng2025uniform, su2025theoretical, karimi2025fs, yue2025oat, cheng2025alpha] models probability transport over discrete token spaces via a continuous-time Markov chain (CTMC) that carries a simple base distribution to the data distribution. Unlike autoregressive decoders that commit to tokens sequentially and accumulate exposure-bias errors, Discrete flow matching supports fully parallel denoising and bidirectional refinement during generation. These properties enable coarse-to-fine planning: beginning with a coarse motion hypothesis, the model increases trajectory fidelity through additional denoising steps, yielding a tunable compute–accuracy trade-off. This flexibility aligns well with autonomous driving, where simple scenes admit rapid approximate plans while complex interactions require higher-precision refinement. Despite these advantages, discrete flow matching remains largely unexplored for VLA policies in end-to-end autonomous driving.
@@
However, a straightforward application of discrete flow matching to VLA model for end-to-end autonomous driving is nontrivial for three reasons. First, training discrete flow matching from scratch is prohibitively data- and compute-intensive, so they are typically initialized from general-purpose autoregressive multimodal VLMs that lack sufficient road-scene competence–from low-level perception and motion forecasting to high-level planning and decision making. We therefore adopt a multi-stage adaptation strategy: starting from a generic VLM backbone (Janus-1.5B [Janus-2025]), we continued conduct pretraining on large-scale road-scene visual question answering (VQA) to strengthen the ability to understand various complex road scenes and vehicle driving patterns, establishing a strong domain prior comparable to autoregressive VLA baselines. Second, standard text token embeddings are ill-suited to high-precision numerical regression because they weakly encode metric relationships. We introduce a metric-aligned numerical tokenizer that discretizes continuous scalars into a shared codebook and learns embeddings with a triplet-margin ranking objective so that latent distances reflect underlying scalar differences. This structured token space enables stable coarse-to-fine and slow–fast trajectory refinement within discrete flow matching, providing a controllable compute–accuracy trade-off. Finally, supervised likelihood-based flow training aligns the model with expert trajectories but does not explicitly enforce safety, ego-progress, and comfort in closed-loop control. We incorporate a Group Relative Policy Optimization (GRPO) based alignment objective with a composite reward that integrates safety penalties and performance goals, improving the safety–progress–comfort profile while preserving the model’s parallel generation capabilities.
@@ 2 Related Work
VLMs in Autonomous Driving. Autoregressive VLMs  [EMMA-2024, OpenEMMA-2025, FutureSightDrive-2025, zhou2025autovla, tian2024drivevlm, jiang2024senna] formulate driving as a sequential language modeling problem, where each token corresponds to a trajectory point, control command, or reasoning step. Representative works such as EMMA [EMMA-2024], OpenEMMA [OpenEMMA-2025], FutureSightDrive [FutureSightDrive-2025] and AutoVLA [zhou2025autovla] leverage chain-of-thought reasoning and external memory modules to enhance interpretability and decision transparency. Despite their strong causal modeling capability, autoregressive architectures suffer from slow autoregressive decoding and limited parallelism, as future actions must be generated step-by-step. Diffusion-based methods, including DiffusionDrive [DiffusionDrive-2025], ViLaD [cui2025vilad] and DiffVLA [jiang2025diffvla] treat planning as a denoising process that gradually refines latent trajectory representations. These models enable parallel sampling but often lack explicit reasoning interpretability. In this paper, we explore a new promising paradigm for end-to-end autonomous driving, namely discrete flow matching.
@@
Discrete Diffusion in LLMs and VLMs. Recent progress in discrete generative modeling has led to the emergence of discrete diffusion LLMs [nie2025llada, ye2025dream] and discrete diffusion VLMs [nie2025large, yu2025dimple, you2025llada-v, li2025lavida, yang2025mmada], which extend diffusion processes to tokenized sequences. This direction originates from D3PM [austin2021-d3pm], which formulated diffusion as a discrete Markov process over categorical variables. Recently, LLaDA [nie2025llada] trains an 8B-parameter model from scratch, reaching LLaMA-3 [grattafiori2024llama-3] performance with bidirectional reasoning and robustness. DREAM-7B [ye2025dream] further enhances diffusion-based reasoning via iterative refinement and arbitrary-order generation. Meanwhile, discrete flow matching [gat2024discrete, lipman2024flow, shaul2024flow] generalizes discrete diffusion via continuous-time probability paths and learnable velocity fields. By unifying diffusion and flow-based generation under a single probabilistic framework, it enables parallel, bidirectional, and efficient sampling. FUDOKI [wang2025fudoki] extends this framework to multimodal reasoning and generation, demonstrating unified, non-autoregressive modeling across modalities. In this paper, we apply discrete flow matching to VLA for autonomous driving and explore its inherent nature of parallel generation and coarse-to-fine controllability.
@@
Reinforcement Learning in VLA. Building upon the success of DeepSeek-R1 [guo2025deepseek], GRPO has been further extended to autonomous driving domains. In particular, AlphaDrive [jiang2025alphadrive] pioneers the integration of GRPO-based reinforcement learning with planning-centric reasoning in autonomous driving, achieving notable improvements in both decision-making performance and training efficiency. TrajHF [li2025finetuning] further combines diffusion-based multimodal planners with reinforcement learning from human feedback, enabling safe and personalized trajectory generation aligned with diverse human driving styles. More recently, AutoVLA [zhou2025autovla] incorporates GRPO into vision-language-action models, extending reinforcement learning to end-to-end multimodal reasoning and low-level planning. To the best of our knowledge, this work presents the first exploration of GRPO within discrete flow matching for autonomous driving VLA. Furthermore, we explicitly incorporate safety alignment objectives, extending beyond conventional likelihood-based training to enhance reliability in autonomous driving contexts.
@@ 3.2 WAM-Flow Architecture
Model Architecture. We adapt a Janus-1.5B multimodal backbone to the discrete flow matching generation paradigm for vision–language–action planning. Images are resized with preserved aspect ratio, zero-padded to 384×384, and encoded by SigLIP [SigLIP-2023] into 576 visual tokens; a lightweight MLP aligns these features to the 2048-dimensional Janus text-token space. On the language side, we extend the Janus tokenizer by 20,001 numerically grounded tokens to represent input ego-state numbers and output waypoint coordinates, yielding a 122,401-word vocabulary. Training data are formatted with a fixed QA-style prompt that integrates navigation commands, ego-state (position, heading, velocity, acceleration), and the target waypoint sequence for the next 4 seconds (8 waypoints). For the decoder, the original Janus text head is expanded to the enlarged vocabulary and used to predict action tokens under the discrete flow matching objective.
@@ 3.4 Training and Inference
Autoregressive-to-Flow Training. Figure 2 outlines a four-stage curriculum. First, we randomly initialize the numerical embeddings and freeze the VLA backbone. We train the numerical embeddings together with the language-model head on the 668K nuPlan dataset for 4 epochs, using flow-matching loss $\mathcal{L}_{\mathrm{CE}}$ (Equation 8) and triplet-margin ranking loss $\mathcal{L}_{\mathrm{num}}$ (Equation 5). Second, we enhance the perception of driving scenes by pretraining VLA using $\mathcal{L}_{\mathrm{CE}}$. This stage trains 3 epochs on 6.5M VQA, including general multimodal VQA (3.4M) from LLaVA-v1.5 [LLaVA] and large-scale driving-specific VQA (3.1M) from RecogDrive [RecogDrive-2025], which enhances perceptual grounding and driving-specific causal reasoning. Third, we supervised fine-tuning of the VLA backbone for only 2 epochs on the nuPlan dataset with $\mathcal{L}_{\mathrm{CE}}$. After supervised flow training, we perform reinforcement learning with simulator feedback by maximizing the GRPO objective (Equation 10) with KL regularization toward the supervised reference to optimize our VLA model for 0.5 epoch on 103k NAVSIM dataset. We set the weight for $\mathrm{EP},\mathrm{TTC}$ and $\mathrm{C}$ in our reward to 5:5:2.
@@ 4.1 Experimental Setup
Metrics. We evaluate our method on the closed-loop NAVSIM-v1 [NAVSIM-v1] and v2 [NAVSIM-v2] benchmarks. The primary metric for NAVSIM-v1 is the Predictive Driver Model Score (PDMS), a composite measure integrating five key components: No-Collision rate (NC), Drivable Area Compliance (DAC), Time-to-Collision within bound (TTC), Comfort, and Ego Progress (EP). The more comprehensive NAVSIM-v2 benchmark employs the Extended Predictive Driver Model Score (EPDMS), which incorporates nine sub-metrics—NC, DAC, Driving Direction Compliance (DDC), Traffic Light Compliance (TLC), EP, TTC, Lane Keeping (LK), History Comfort (HC), and Extended Comfort (EC)– to provide a holistic assessment of driving performance, safety, and rule adherence. All results are obtained from closed-loop simulations on the official public test splits.
@@ 5 Conclusion
This appendix provides additional experimental results and implementation details to complement the main paper. Specifically, Section A presents extended evaluations on the nuScenes [nuScenes-dataset-2020] datasets, along with additional qualitative experiments on NAVSIM. Section B provides pseudocode for the training and inference stages, respectively. Section C elaborates on the evaluation metrics, and Section D discusses implementation specifics. Finally, Section E discuss the limitation and future work.
@@ Appendix A Additional Experiments
We evaluate our method on the nuScenes dataset [nuScenes-dataset-2020] following the NAVSIM benchmark perspective [NAVSIM-v1, NAVSIM-v2], which focuses on collision rate as the primary metric. This emphasis stems from the established finding in NAVSIM that open-loop L2 distance exhibits negligible correlation with closed-loop performance. As shown in Table 7, our method achieves an average collision rate of 0.12% under ST-P3 metrics, matching the performance of the best non-VLA model (UniAD). More notably, under the more comprehensive UniAD metrics, WAM-Flow sets a new state-of-the-art with the lowest average collision rate (0.23%) among all evaluated VLA methods. The model also demonstrates superior short-term safety, achieving a perfect 0.00% collision rate at the 1-second horizon.
@@
PreWorld [Preworld-2025] - - - - - - 0.19 0.57 2.65 1.14
ST-P3 [ST-P3-2022] - - 0.23 0.62 1.27 0.71 - - - -
Ego-MLP [Ego-MLP-2024] - - 0.21 0.35 0.58 0.38 - - - -
InsightDrive [InsightDrive-2025] - - 0.09 0.10 0.27 0.15 0.08 0.15 0.84 0.36
VAD-v2 [VADv2-2024] - - 0.07 0.10 0.24 0.14 - - - -
UniAD [UniAD-2023] - - 0.04 0.08 0.23 0.12 0.05 0.17 0.71 0.31
@@
Epona [Epona-2025] AR + Diff. DiT-2.5B [DiT-2023] 0.05 0.22 0.85 0.96 - - - -
OmniDrive [OminiDrive-2025] AR LLaVA-7B [LLaVA] 0.04 0.46 2.32 0.94 - - - -
DriveVLM [tian2024drivevlm] AR Qwen2-VL-7B [wang2024qwen2] 0.10 0.22 0.45 0.27 - - - -
GPT-Driver [GPT-Driver-2023] AR GPT-4 [GPT-4-2023] 0.04 0.12 0.36 0.17 0.07 0.15 1.10 0.44
AutoVLA [zhou2025autovla] AR Qwen2.5-3B [Qwen-25] 0.13 0.18 0.28 0.20 0.14 0.25 0.53 0.31
DME-Driver [DME-Driver-2025] AR LLaVA-7B [LLaVA] - - - - 0.05 0.28 0.55 0.29
Ours DFM Janus-1.5B [Janus-2025] 0.04 0.10 0.23 0.12 0.00 0.10 0.60 0.23
Table 7: End-to-end motion planning performance on the nuScenes [nuScenes-dataset-2020] dataset. We sort previous methods according to the average collision rate. Abbreviation: Diff.(Diffusion), AR (autoregressive), DFM (discrete flow matching).